Train cars are coupled together by being bumped in
Answer #1

  This is an inelastic collision so only momentum is conserved. Initial momentum of the system of cars, Pi, is

Pi = m1*v1 + m2*v2 = 170000*0.3 kg-m/s + 115000*(-0.12) kg-m/s = 3.72x10^4 kg-m/s

Now the final momentum Pf = (m1+m2)*v3 so

Pi =Pf ---> Pi = (m1+m2)*v3 ---> v3 = Pi/(m1+m2) = 3.72x10^4 (kg-ms/)(170,000+115,000 kg)

v3 = 0.130 m/s in the direction that m1 was travelling in